Capt4d:

The sum of three consecutive numbers is 84. What is the largest of these numbers?

UzerMate:

you got choices?

Hoodmemes:

Do u know what "Constructive numbers" are?

Capt4d:

26 27 28 29

XioGonz:

Sum up three numbers at a time, once you do that and your answer outcome is 84, you have to see which of the numbers you added is the largest.

Hoodmemes:

Honestly,to get your answer u would first do this \[84 \div 3 = 28\] And as we all know \[28+28+28=82\] So u would add 1 to one of the 28's and minus 1 from 1 of the 28's like dis \[ 1 - 28=(blank) and ,28-1=(blank) \] The answer u get for each will be one of your answer to which is two of the largest,and 28 would be one too

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes Honestly,to get your answer u would first do this \[84 \div 3 = 28\] And as we all know \[28+28+28=82\] So u would add 1 to one of the 28's and minus 1 from 1 of the 28's like dis \[ 1 - 28=(blank) and ,28-1=(blank) \] The answer u get for each will be one of your answer to which is two of the largest,and 28 would be one too \(\color{#0cbb34}{\text{End of Quote}}\) i mean 1+28 oof